OpenStudy (countrygirl1431):

Please Help Me!!!!! Algebra 2 Place the indicated product in the proper location on the grid. (3m 3 - 1/2 y )(3m 3 - 1/2 y )

OpenStudy (dayakar):

do u know x*x =

OpenStudy (countrygirl1431):

no i dont know anything about this

OpenStudy (dayakar):

do u know the identity \[\left( a-b \right)^{2}=\]

OpenStudy (countrygirl1431):

no

OpenStudy (dayakar):

can u multiply (a-b)(a-b)=

OpenStudy (countrygirl1431):

i guess

OpenStudy (dayakar):

very good try it

OpenStudy (dayakar):

what happens

OpenStudy (countrygirl1431):

idk im completly confused

OpenStudy (dayakar):